# The Geometric View of Theories

### v1

@@ B.1 A wave-function is a CFT partition function: models “one level up”
and the $h$’s are linear perturbations of the spatial metric. Here $z_{1},\ldots,z_{n}$ are “holographic” coordinates at timelike future infinity, denoted collectively by $z$, that are integrated over. $\{ij\}$ denotes the (symmetrized) collections of indices $(i_{1},\ldots,i_{n})$ and $(j_{1},\ldots,j_{n})$, and the Einstein summation convention for repeated indices is used. For the technical details, see Chakraborty et al. (2024: pp. 20-21). The main point is that, near the Euclidean vacuum, this allows us to write the solution of the WDW equation in a similar way to the above expression, i.e. as an expansion in powers of Newton’s constant.
@@ B.2 A normed vector space
The above set of wave-functions $\Psi[g]$ was given in a perturbative expansion, around the Euclidean vacuum. This space of solutions comes with a natural norm. This norm is obtained by integrating $|Z[g]|^{2}$ over the spatial metrics, and dividing by the volume of its symmetry group, namely the diffeomorphism and Weyl groups. [^65] Technically, this is done using the Faddeev-Popov procedure: see Chakraborty et al. (2023: pp. 9-12). The explicit expression for this norm can be found in Chakraborty et al. (2023: pp. 4, 10). It is important to note that it is valid beyond the limit $\kappa\rightarrow 0$. In particular, its $\kappa\rightarrow 0$ limit reproduces the earlier norm written by Higuchi for quantum field theory in de Sitter space, in terms of de Sitter-invariant states. However, the Chakraborty et al. (2023) norm goes beyond this limit. [^66] There are some interesting subtleties about this limit that I cannot address here, and which signal significant differences between quantum gravity and quantum field theory. For a discussion, see Chakraborty et al. (2023: p. 21).
The main upshot of this discussion is that it gives a vivid illustration, for a cosmologically relevant treatment of quantum cosmology, of the two main aspects of the geometric view. Namely, the partition function of a CFT is “merely” a wave-function, i.e. one element of the state-space of de Sitter quantum gravity. Thus what we normally think of as being a theory, is here rightly thought of as a model. Second, there is further structure on the set of CFTs, i.e. on the set of models: namely, it is a normed vector space. While this latter propery is part of the definition of a state-space, it is not in general to be expected of the set of CFTs. It is the special set of (non-unitary) CFTs that make up the theory of quantum cosmology, that can be given this vector space and normed structure. [^67] It is only ‘normed’ structure and not an inner product, because these papers do not work out the inner products between different vectors. Hence my calling it a state-space rather than a Hilbert space.

### v2

@@ B.1 A wave-function is a CFT partition function: models “one level up”
and the $h$’s are linear perturbations of the spatial metric. Here $z_{1},\ldots,z_{n}$ are “holographic” coordinates at timelike future infinity, denoted collectively by $z$, that are integrated over. $\{ij\}$ denotes the (symmetrized) collections of indices $(i_{1},\ldots,i_{n})$ and $(j_{1},\ldots,j_{n})$, and the Einstein summation convention for repeated indices is used. For the technical details, see Chakraborty et al. (2024: pp. 20–21). The main point is that, near the Euclidean vacuum, this allows us to write the solution of the WDW equation in a similar way to the above expression, i.e. as an expansion in powers of Newton’s constant.
@@ B.2 A normed vector space
The above set of wave-functions $\Psi[g]$ was given in a perturbative expansion, around the Euclidean vacuum. This space of solutions comes with a natural norm. This norm is obtained by integrating $|Z[g]|^{2}$ over the spatial metrics, and dividing by the volume of its symmetry group, namely the diffeomorphism and Weyl groups. [^71] Technically, this is done using the Faddeev-Popov procedure: see Chakraborty et al. (2023: pp. 9–12). The explicit expression for this norm can be found in Chakraborty et al. (2023: pp. 4, 10). It is important to note that it is valid beyond the limit $\kappa\rightarrow 0$. In particular, its $\kappa\rightarrow 0$ limit reproduces the earlier norm written by Higuchi for quantum field theory in de Sitter space, in terms of de Sitter-invariant states. However, the Chakraborty et al. (2023) norm goes beyond this limit. [^72] There are some interesting subtleties about this limit that I cannot address here, and which signal significant differences between quantum gravity and quantum field theory. For a discussion, see Chakraborty et al. (2023: p. 21).
The main upshot of this discussion is that it gives a vivid illustration, for a cosmologically relevant treatment of quantum cosmology, of the two main aspects of the geometric view. Namely, the partition function of a CFT is “merely” a wave-function, i.e. one element of the state-space of de Sitter quantum gravity. Thus what we normally think of as being a theory, is here rightly thought of as a model. Second, there is further structure on the set of CFTs, i.e. on the set of models: namely, it is a normed vector space. While this latter propery is part of the definition of a state-space, it is not in general to be expected of the set of CFTs. It is the special set of (non-unitary) CFTs that make up the theory of quantum cosmology, that can be given this vector space and normed structure. [^73] It is only ‘normed’ structure and not an inner product, because these papers do not work out the inner products between different vectors. Hence my calling it a state-space rather than a Hilbert space. For a general discussion of probability measures in the context of classical cosmology, see Curiel (2015).